September 5, 2013 FDC meeting
- Installation preparations 
- Engineering (Bill)
- Electronics (Chris, Nick)
- Tests at EEL126 
Participants: Bill, Dave, Chris, Nick, Simon, Vlad, Eugene, and Lubomir.
- We discussed briefly the status in the Hall with respect to what is needed for the FDC installation. The electronics racks are on one side of the platform but not installed yet (Dave), rather just used as a template to mark the positions of the cable opening. After BCAL is installed the BCAL platform will be removed; for the cabling of the bottom of the FDC we will use something temporary there. At the same time the transformers should be installed there, but not clear how much space they will occupy.
- Gas system: nothing is in the Hall yet. There are penetration between the gas room and the Hall, but not yet to the gas bottles outside. The plan is to use plastic tubes from the gas room through the Hall to the detector. We want to check if using such a long plastic tubes will increase the Oxygen contamination. We can use the Oxygen sensor right now connected to the spare package. In the next weeks when Beni is back, we want to discuss the gas control with Narciso and Dave.
- Bill will check if the chilled water is readily available, and how and when the dry air system will be set-up.
- Bill attended a kind of QA readiness review for the FDC cabling. He explained the need to remove the four extra rails for the mesh; Tim agreed. One of the suggestion was to consult with Dr.Chandler about the ergonomics of the cabling (see the second picture attached above).
- The strips to connect the bearings of the four packages can be used as a ground connection and no additional ground cable is needed for that; Fernando agreed. We decided to have the strips plated for better surface conductivity. In addition Bill will use washers with teeth to connect to the bearing plates. It is not clear yet if one can use also the mesh as a ground connection. At the same time it may serve as a Faraday cage for the CDC!
- Bill and Dave will figure out when there will be a tech available to work on the skin: making slots on the g10 and cutting the plastic holders. We will need also Casey to install the tubing between the cooling manifolds.
- Bill wants to test the new mesh support in the ESB. Only after that we will remove the four extra rails together with their holders on the installation carts.
- Chris and Nick fixed the two HV connectors on package #2 that had several wires disconnected. Vlad tested the package with HV and all the sectors had signals.
Tests at EEL126
- Vlad finished testing the FDC packages with HV. He applies HV package by package and keep it for at least 24 hours. He checks two wire discriminator cards (15V threshold) in the middle of each cell. Each middle card has wires connected to two HV sectors. Thus, he checked that the positive HV is OK on all the sectors but there's no way to say if the negative is there too. Out of ~1k channels tested he found two noisy wires. Vald will keep the HV for longer time to see if the noise will disappear. It might be also due to the very low gas flow, 50ccpm per package to be compared to the nominal 250ccpm.